# LTGC: Long-tail Recognition via Leveraging LLMs-driven Generated Content

### v3

@@ Abstract
Long-tail recognition is challenging because it requires the model to learn good representations from tail categories and address imbalances across all categories. In this paper, we propose a novel generative and fine-tuning framework, LTGC, to handle long-tail recognition via leveraging generated content. Firstly, inspired by the rich implicit knowledge in large-scale models (e.g., large language models, LLMs), LTGC leverages the power of these models to parse and reason over the original tail data to produce diverse tail-class content. We then propose several novel designs for LTGC to ensure the quality of the generated data and to efficiently fine-tune the model using both the generated and original data. The visualization demonstrates the effectiveness of the generation module in LTGC, which produces accurate and diverse tail data. Additionally, the experimental results demonstrate that our LTGC outperforms existing state-of-the-art methods on popular long-tailed benchmarks.
@@ 1 Introduction
In the real world, data often exhibits a long-tailed distribution, posing significant challenges for computer vision recognition [zhang2023deep, 46]. These challenges include (1) Class Imbalance: within the dataset, some classes (termed ”head” classes) are abundantly represented, while others (termed ”tail” classes) have few samples. This imbalanced distribution during training may cause the model to focus more on the head classes, neglecting the tail classes [44]. (2) Tail Data Scarcity: Data scarcity refers to tail classes having an extremely limited number of samples, which lack diversity and are insufficient to effectively train a model [24, 8]. It prevents a model’s ability to learn the feature invariant, which is necessary to recognize these categories correctly [35].
3 Refer to caption
@@
To address these challenges, numerous approaches have emerged within the field of long-tailed recognition, such as re-sampling [5, 19], loss re-weighting [23, 26, 12, 6, 50, 1, 29, 27, 41], ensemble learning [49, 44, 57], decoupling [20], and contrastive learning [42, 61, 11, 25]. Their primary objective is to balance the decision boundaries and parameter weights of the model to improve long-tail recognition. However, these methods encounter bottlenecks due to the scarcity and limited diversity features of tail-class data [24]. To obtain more diverse representations for tail classes, two types of methods appear: Some methods increase tail-class diversity through data augmentation [9, 24, mixup, zhao2022mixpro]. Others increase tail-class diversity by transferring features from related classes [8] or large pre-trained models (e.g., CLIP [32]). These transfer learning methods [14, 28, 36] based on CLIP have recently shown great potential in boosting long-tail recognition. The LPT [14] fine-tunes pre-trained models to adapt to target datasets. The VL-LTR [36] leverages text features from CLIP to augment the learning of image features. However, these methods find it hard to obtain the correct and desired knowledge for tail categories. For example, in the real-world data iNauralist 2018 [17], the category Aquilegia Pubescens is characterized only by white or pale yellow colors and spurred shapes. If we transfer similar category features or blindly augment semantic information, such as red color semantics, this category will become more confusing.
@@
Recently, Large Language Models (LLMs) (e.g., ChatGPT) and Large Multimodal Models (LMMs) (e.g., GPT-4V (ChatGPT with Vision [52]) and Minigpt4 [60, 7]) due to their wealth of implicit knowledge, have been leveraged for a variety of downstream tasks, such as robot task plans [34], open-set object recognition [30], automated robot learning [45] and various visual reasoning tasks [52, 53, zhu2023llafs]. However, due to the bias of imbalanced training data, LMMs perform poorly on some complex tasks [10]. We also find this dilemma in long-tail recognition as shown in Fig. 1. When we asked LMMs about an image in the category of Trogon rufus: ”What species is in the image? ”, we did not get the expected answer. In contrast, when we ask some easy questions about this image: ”Please describe the image.” or ”Please describe the distinctive features of Trogon Rufus”, we get the desired answer. This suggests that although large models underperform in long-tail recognition (e.g., it is difficult to align complex image features with labels.), they still contain desired implicit knowledge (e.g., correctly describing images and providing the knowledge of species). Inspired by this, we aim to leverage the rich implicit knowledge of several large models to address the challenges of long-tailed data scarcity and perform long-tailed recognition well. Nevertheless, effectively leveraging large models to encounter these challenges is not trivial: (1) The implicit nature of knowledge within large models makes it difficult to extract the desired knowledge that facilitates long-tail recognition tasks. (2) Due to the domain gap between the generated data and the original data [38], effectively using these hybrid data remains an open issue.
@@
To this end, we propose a novel Long-Tail recognition framework via Generated Content, denoted as LTGC, which is illustrated in Fig. 2. Our LTGC aims to leverage the capabilities of large models for generating explicitly diverse content [foo2023aigenerated] tailored to the long-tail classes and incorporate novel designs to enhance long-tail recognition. Firstly, motivated by the wealth of implicit knowledge in large models and the fact that texts are more controllable, we employ them to expand the tail classes. Specifically, to produce more diverse and controllable tail data, we follow the rule of analyzing existing features before generating absent features. To achieve this, LTGC utilizes LMMs to analyze images already in the tail classes to obtain the existing tail-class descriptions list. Then LTGC inputs the existing descriptions list to LLMs to obtain the desired extended tail-class descriptions list for images that are absent from the existing tail classes. Secondly, LTGC utilizes the text-to-image (T2I) model (e.g., DALL-E [33]) to transform these textual descriptions into images. Moreover, inspired by the benefits of chain-of-thought [48], we propose a self-reflection and iterative evaluation module for this process to ensure the diversity and quality of the generated content. Finally, inspired by the advantages of Mixup [54] in merging different image domains, we propose the BalanceMix module to address domain shifts of generated images for the fine-tuning process.
@@ 2 Related Work
Large Models. Recently, several large models have emerged, including large language models (e.g., ChatGPT [52]), large multi-modal models (e.g., GPT-4 (ision) [52], MiniGPT [60]) and large generated models (e.g., DALL-E [33]). These expansive models are repositories of extensive knowledge [foo2023aigenerated] and have undergone scrutiny in diverse applications, such as robot task plans [34], open-set object recognition [30], contextual object detection [53], video generation [22], and few-shot segmentation [zhu2023llafs]. In this work, we leverage the rich implicit knowledge of these large models to handle the long-tail recognition challenge.
@@ 5 Conclusion
[…]
Supplementary Material
A1 Experiments Details
A1.1 Dataset
We conduct the experiments on three popular long-tailed benchmarks, including ImageNet-LT [27], Places-LT [27], and iNaturalist 2018 [17]. We present these datasets in detail below.
ImageNet-LT. The ImageNet-LT [27] is the long-tailed version of the dataset ImageNet-2012 [13]. Overall, ImageNet-LT has 115.8K images from 1000 categories with an imbalanced factor $\beta=1280/5$.
Places-LT is a long-tailed version of the large-scale scene classification dataset Places [59]. It consists of 62.5K images from 365 categories with class cardinality ranging from 5 to 4,980.
iNaturalist 2018. The iNaturalist 2018 [17] is a large-scale dataset for long-tailed visual recognition rooted in real-world scenarios and exhibits a highly imbalanced distribution. It encompasses 437.5K training images and 24.4K validation images spread across 8142 categories. The dataset’s fine-grained nature further intensifies its complexity.
A1.2 More Results on ImageNet-LT and Place-LT
As shown in Tab. A1 and Tab. A2, we present additional experimental results of our proposed LTGC on ImageNet-LT and Place-LT datasets, including results on the ’Many’ and ’Medium’ split sets.
Table A1: Comparison with SOTA methods on ImageNet-LT.
Method Many Medium Few All
CLIP Zero [36] 60.8 59.3 58.6 59.8
CLIP Finetune [36] 74.4 56.9 34.5 60.5
VL-LTR [36] 77.8 67.0 50.8 70.1
LTGC(Ours) 83.0 79.8 70.5 80.6
Table A2: Comparison with SOTA methods on Places-LT.
Method Many Medium Few All
CLIP Zero [36] 37.5 37.5 40.1 38.0
CLIP Finetune [36] 50.8 38.6 22.7 39.7
VL-LTR [36] 54.2 48.5 42.0 50.1
RAC [28] 48.7 48.3 41.8 47.2
LPT [14] 49.3 52.3 46.9 50.1
LTGC(Ours) 55.8 54.2 52.1 54.1
A1.3 Prompts for LMMs comparison.
For the experiments on LMMs (Sec. 4.3), it’s worth noting that providing a labels list significantly improves the model’s performance compared to querying without a list (i.e., directly asking ’What species is in the image?’). The labels list provides prior information to the model, helping it narrow down the decision space within its extensive knowledge base. Therefore, on ImageNet-LT, we used the query template ’Please classify this image. Choose from the following classes: Class 1, Class 2, …, Class Y.’.
However, it is essential to note that the iNaturalist 2018 dataset encompasses a substantial number of classes, totaling 8,141, which is approximately 184 KB of text. Currently, no LLMs (including the most advanced GPT-4) are capable of processing such a lengthy input. Inputting all categories at once would lead to an error due to the limitations in text length. Dividing the categories into multiple inputs would cause the model to forget previous categories, leading to inaccurate judgments based on the most recent inputs. Therefore, for the iNaturalist dataset, we adopted the query template ’What species is in the image?’ for testing purposes. We then reviewed the entire response, considering the LLMs’ classification as correct if it included any vocabulary matching the Ground Truth class, and as incorrect otherwise.
Table A3: Effectiveness of the Self-reflection module. NC: Number-checking. RC: Repetition-checking.
Dataset NC RC Top-1 Accuracy
ImageNet-LT - - 60.5
- 71.7
- 73.2
80.6
iNaturalist 2018 - - 71.6
- 78.4
- 71.8
82.5
A2 Additional Ablation Studies
Here, we conduct additional ablation experiments. Unless otherwise noted, we report the top-1 accuracy averaged over three runs on the ImageNet-LT evaluation protocol.
Effectiveness of the Self-reflection module. The self-reflection module consists of two elements: a number check for images and a repetition check. We separately investigated these two checks, and the results are shown in Tab. A3. When only performing the quality check, LLMs generate highly repetitive descriptions, leading to a decrease in textual diversity and, consequently, a decline in image diversity. When only performing the repetitiveness check, the LLMs are posed the two questions mentioned in Section 3.1 only once. This results in a limited number of generated samples, thereby leading to limited performance gains. As shown in Tab. A3, our proposed method incorporates both checks and consistently outperforms all variants. This attests to (1) the effectiveness of the repetitiveness design, enabling LLMs to generate comprehensive text descriptions, and (2) the efficacy of the self-reflection design, allowing LLMs to increase the number of samples and simulate a more balanced set of classes.
Impact of using different versions of ChatGPT of LLM. Besides, we also test using different versions of ChatGPT with different capabilities, including GPT 3.0, GPT-3.5 Turbo 16K, and the latest GPT-4 Turbo 128K. As shown in Tab. A4, our framework with different versions of GPT used can achieve different results. This shows that the performance of our framework is affected by the version of ChatGPT.
Table A4: Evaluation on using different versions of ChatGPT.
Version ImageNet-LT iNaturalist
GPT 3.0 72.0 74.1
GPT 3.5 76.9 79.3
GPT 4.0 80.6 82.5
The effectiveness of different maximum image numbers for LLM’s self-reflection module. In our framework, the maximum number of generated and original images for each class $y$, denoted as $K_{y}$. For LLM’s self-reflection module, the $K_{y}$ is set to 100 for iNaturalist 2018, 300 for ImageNet-LT, and 800 for Place-LT, respectively. We explore alternative caps for the maximum image number and show the findings in Fig. A1. The results indicate that for each dataset when the cap is set below 100, 300, and 800, respectively, there’s a noticeable improvement in our framework’s performance as the limit on generative images increases. This might be because, by setting the maximum number of images to be a larger number, LLM can better cover the diverse distinctive features and backgrounds (scenes). Moreover, it’s observed that increasing the maximum image number beyond 100, 300, and 800 does not lead to further improvements in performance. However, surpassing these maximum numbers of 100, 300, and 800 doesn’t yield further performance gains. Therefore, balancing performance and efficiency, we determined these respective limits as the optimal settings for our framework.
Figure A1: The effectiveness of different maximum image numbers for LLM’s self-reflection module.
Impact of the maximum number of cycle times for iterative evaluation module. In our framework’s iterative evaluation module, we initially cap the cycle count at three. To explore the impact of varying this limit, we experimented with different maximum cycle thresholds and presented the outcomes in Table A5. The results indicate an enhancement in our framework’s performance as we expand the maximum cycle count for iterative evaluation. This improvement is likely due to the more thorough refinement of generated images achieved with a higher cycle limit. However, the gains in performance diminish when the cycle count exceeds three, leading us to establish three as the optimal maximum for the iterative evaluation module.
Table A5: Evaluation on the maximum number of cycle times for iterative evaluation.
Maximum number of cycle time ImageNet-LT
1 71.5
2 78.1
3 80.6
4 79.9
Check the reasonability of the generated descriptions. In our framework, we use ChatGPT to generate text descriptions for the evaluated classes, and we further design a iterative evaluation module in which we guide ChatGPT to modify its generated descriptions. Here, we perform a check w.r.t. the reasonability of the generated descriptions before and after passing into the iterative evaluation module. Specifically, we find that, before passing into the iterative evaluation module, 4% of descriptions are checked to be unreasonable, but 0.2% of the descriptions output by the module is checked to be unreasonable. This shows that ChatGPT has small probability of generating unreasonable descriptions, while our iterative evaluation module can further mitigate this problem. The above check is done by inviting 3 volunteers and passing the same 1000 descriptions to them. The 3 volunteers first make decisions independently and then discuss disagreed decisions.
Figure A2: Visualization about the iterative evaluation module. The images in (a) do not align most closely with their intended classes until the third cycle, whereas the images in (b) achieve their closest alignment with the intended classes in the second cycle. The classes of Tamias Amoenus and Lepus Townsendii are derived from the iNaturalist dataset, while Swimming Pool and Triumphal Arch are from the Place-LT and ImageNet-LT datasets, respectively.
Preparation time. In our framework, before entering the fine-tuning process, we first prepare the generated images by extending the descriptions of tail classes and then generate diverse images for each class, filtering and regenerating the low-quality images. We here report the preparation time of our framework. For generating images, the speed of image generation is limited by DALL-E’s API, about 50 images/minute [33]. For generating descriptions, the speed is about 300 items/minute. Note that the entire preparation process of our framework can be automated by a script with multi-threaded accelerated generation.
The effectiveness of fine-tuning with LoRA. In the fine-tuning process, we fine-tune the visual encoder of the CLIP model using Low-Rank Adaptation (LoRA) [18], LoRA layers are typically added to the Transformer layers of the CLIP model. The main idea of LoRA is to make low-rank modifications to existing weight matrices, rather than updating all parameters directly. This method is often used in large-scale models to achieve effective and parameter-efficient fine-tuning.
Each Transformer layer in the CLIP model consists of two main components: the Multi-Head Self-Attention and the Feedforward Neural Network. LoRA can be applied to the weight matrices of both these components.
Specifically, in the Multi-Head Self-Attention, LoRA can be applied to the Query, Key, and Value matrices. For the Feedforward Network, LoRA can be integrated into the first linear transformation, which is right before the ReLU activation function.
Moreover, other competitive methods of parameter fine-tuning have emerged in recent years, such as VQT [39]. In Tab. A6 we compare LoRA with other fine-tuning methods such as Linear-probing [32] and VQT [39]. We observe that LoRA outperforms other fine-tuning methods, demonstrating the effectiveness of our fine-tuning with LoRA
Table A6: The effectiveness of fine-tuning with LoRA.
Method ImageNet-LT iNaturalist
Linear-probing [32] 71.4 66.3
VQT [39] 77.3 78.1
LoRA (Ours) 80.6 82.5
Figure A3: Domain gap between generated images and original images.
A3 Additional Visualizations
Visualisation of the iterative evaluation module. In this work, we propose an iterative evaluation module for our framework. This module aims to enhance the quality of generated images, particularly those that are not highly accurate. Fig A2 illustrates the refinement process carried out by our iterative evaluation module. As depicted, this module effectively improves images that initially do not align well with their intended classes, ensuring they more accurately represent these classes. This highlights the efficiency and utility of our iterative evaluation module in image refinement.
Domain gap between generated images and original images. Some previous methods have shown that there is a domain gap [38] between the generated data and the original data, to better illustrate the domain gap we visualize the t-SNE [40] results for the part of the original data and the part of generated data. In Fig. A3, we observe that there is a domain gap between the generated data and the original data. This motivated us to propose the BalanceMix method and significantly helped the model fine-tuning to improve the recognition results.
A4 Licenses
Large model licenses. We use ChatGPT and DALL-E by following the terms of using the services of OpenAI. We use CLIP by following the MIT License.

### v4

@@ Abstract
Long-tail recognition is challenging because it requires the model to learn good representations from tail categories and address imbalances across all categories. In this paper, we propose a novel generative and fine-tuning framework, LTGC, to handle long-tail recognition via leveraging generated content. Firstly, inspired by the rich implicit knowledge in large-scale models (e.g., large language models, LLMs), LTGC leverages the power of these models to parse and reason over the original tail data to produce diverse tail-class content. We then propose several novel designs for LTGC to ensure the quality of the generated data and to efficiently fine-tune the model using both the generated and original data. The visualization demonstrates the effectiveness of the generation module in LTGC, which produces accurate and diverse tail data. Additionally, the experimental results demonstrate that our LTGC outperforms existing state-of-the-art methods on popular long-tailed benchmarks. Project Link.
@@ 1 Introduction
In the real world, data often exhibits a long-tailed distribution, posing significant challenges for computer vision recognition [57, 47]. These challenges include (1) Class Imbalance: within the dataset, some classes (termed ”head” classes) are abundantly represented, while others (termed ”tail” classes) have few samples. This imbalanced distribution during training may cause the model to focus more on the head classes, neglecting the tail classes [45]. (2) Tail Data Scarcity: Data scarcity refers to tail classes having an extremely limited number of samples, which lack diversity and are insufficient to effectively train a model [25, 8]. It prevents a model’s ability to learn the feature invariant, which is necessary to recognize these categories correctly [36].
@@
To address these challenges, numerous approaches have emerged within the field of long-tailed recognition, such as re-sampling [5, 20], loss re-weighting [24, 27, 12, 6, 51, 1, 30, 28, 42], ensemble learning [50, 45, 59], decoupling [21], and contrastive learning [43, 63, 11, 26]. Their primary objective is to balance the decision boundaries and parameter weights of the model to improve long-tail recognition. However, these methods encounter bottlenecks due to the scarcity and limited diversity features of tail-class data [25]. To obtain more diverse representations for tail classes, two types of methods appear: Some methods increase tail-class diversity through data augmentation [9, 25, 55, 58]. Others increase tail-class diversity by transferring features from related classes [8] or large pre-trained models (e.g., CLIP [33]). These transfer learning methods [14, 29, 37] based on CLIP have recently shown great potential in boosting long-tail recognition. The LPT [14] fine-tunes pre-trained models to adapt to target datasets. The VL-LTR [37] leverages text features from CLIP to augment the learning of image features. However, these methods find it hard to obtain the correct and desired knowledge for tail categories. For example, in the real-world data iNauralist 2018 [18], the category Aquilegia Pubescens is characterized only by white or pale yellow colors and spurred shapes. If we transfer similar category features or blindly augment semantic information, such as red color semantics, this category will become more confusing.
@@
Recently, Large Language Models (LLMs) (e.g., ChatGPT) and Large Multimodal Models (LMMs) (e.g., GPT-4V (ChatGPT with Vision [53]) and Minigpt4 [62, 7]) due to their wealth of implicit knowledge, have been leveraged for a variety of downstream tasks, such as robot task plans [35], open-set object recognition [31], automated robot learning [46] and various visual reasoning tasks [53, 54, 64]. However, due to the bias of imbalanced training data, LMMs perform poorly on some complex tasks [10]. We also find this dilemma in long-tail recognition as shown in Fig. 1. When we asked LMMs about an image in the category of Trogon rufus: ”What species is in the image? ”, we did not get the expected answer. In contrast, when we ask some easy questions about this image: ”Please describe the image.” or ”Please describe the distinctive features of Trogon Rufus”, we get the desired answer. This suggests that although large models underperform in long-tail recognition (e.g., it is difficult to align complex image features with labels.), they still contain desired implicit knowledge (e.g., correctly describing images and providing the knowledge of species). Inspired by this, we aim to leverage the rich implicit knowledge of several large models to address the challenges of long-tailed data scarcity and perform long-tailed recognition well. Nevertheless, effectively leveraging large models to encounter these challenges is not trivial: (1) The implicit nature of knowledge within large models makes it difficult to extract the desired knowledge that facilitates long-tail recognition tasks. (2) Due to the domain gap between the generated data and the original data [39], effectively using these hybrid data remains an open issue.
@@
To this end, we propose a novel Long-Tail recognition framework via Generated Content, denoted as LTGC, which is illustrated in Fig. 2. Our LTGC aims to leverage the capabilities of large models for generating explicitly diverse content [16] tailored to the long-tail classes and incorporate novel designs to enhance long-tail recognition. Firstly, motivated by the wealth of implicit knowledge in large models and the fact that texts are more controllable, we employ them to expand the tail classes. Specifically, to produce more diverse and controllable tail data, we follow the rule of analyzing existing features before generating absent features. To achieve this, LTGC utilizes LMMs to analyze images already in the tail classes to obtain the existing tail-class descriptions list. Then LTGC inputs the existing descriptions list to LLMs to obtain the desired extended tail-class descriptions list for images that are absent from the existing tail classes. Secondly, LTGC utilizes the text-to-image (T2I) model (e.g., DALL-E [34]) to transform these textual descriptions into images. Moreover, inspired by the benefits of chain-of-thought [49], we propose a self-reflection and iterative evaluation module for this process to ensure the diversity and quality of the generated content. Finally, inspired by the advantages of Mixup [55] in merging different image domains, we propose the BalanceMix module to address domain shifts of generated images for the fine-tuning process.
@@ 2 Related Work
Large Models. Recently, several large models have emerged, including large language models (e.g., ChatGPT [53]), large multi-modal models (e.g., GPT-4 (ision) [53], MiniGPT [62]) and large generated models (e.g., DALL-E [34]). These expansive models are repositories of extensive knowledge [16] and have undergone scrutiny in diverse applications, such as robot task plans [35], open-set object recognition [31], contextual object detection [54], video generation [23], and few-shot segmentation [64]. In this work, we leverage the rich implicit knowledge of these large models to handle the long-tail recognition challenge.